OpenStudy (nwolf):

-10<=3x-3<=-2 I was able to get -7/3 and 1/3 is the correct

OpenStudy (anonymous):

\[\Large -10\leq 3x-3\leq -2\] \[\Large -10+3\leq 3x\leq -2+3\] \[\Large \frac{-7}{3}\;\;\leq x\;\;\leq \frac11\] \[\Large x\in \left[\frac{-7}{3},1\right]\]

OpenStudy (anonymous):

sorry I made a mistake.. I'll correct that.. hold on ...

OpenStudy (anonymous):

\[\Large \frac{-7}{3}\leq x\leq \frac13\] \[\LARGE x\in\left[\frac{-7}{3},\frac{1}{3}\right]\]